# Comparative Study of Recurrent Neural Networks for Virtual Analog Audio Effects Modeling

### v2

@@ 5 Conclusion
LSTM-based networks excel in replicating distortion and equalizer effects, showing their ability to effectively encode the audio dependencies enough for these specific effects. On the other hand, LSTM when designed in an encoder-decoder fashion, and in particular, SSMs outperform other models when emulating saturation and compression. In the case of compression, they demonstrate their advantage in accurately capturing the signal’s history. Conversely, LRUs produce less favorable outcomes, leading to noticeable audio artifacts. Generally, SSMs have proved to be an appropriate candidate technique also in VA, although the gap with the other models is not always so pronounced.
In this study, we did not carry out architecture- and effect-specific optimizations, which can definitely boost the accuracy, such as hyper-parameters tuning and changing the size of the various layers. We limited computational complexity and input-output latency and focused the study on models that can be theoretically implemented in consumer-grade devices for live audio applications.
In future work, conditioning techniques and performances should be further investigated. The current conditioning blocks rely on affine transformation, but devices such as distortion or saturation could benefit from more complex transformation to take into account the possible nonlinear influence of the parameters.

### v3

@@ 5 Conclusion
LSTM-based networks excel in replicating distortion and equalizer effects, showing their ability to encode the audio dependencies enough for these specific effects effectively. On the other hand, LSTM, when designed in an encoder-decoder fashion, LRUs and, in particular, SSMs, outperform other models when emulating saturation and compression. In the case of compression, they demonstrate their advantage in accurately capturing the signal’s history. Generally, SSMs have also proved to be an appropriate candidate technique in VA, although the gap with the other models is not always so pronounced.
In this study, we did not carry out architecture—and effect-specific optimizations, which can definitely boost the accuracy, such as tuning hyperparameters and changing the size of the various layers. We limited computational complexity and input-output latency and focused the study on models that can be theoretically implemented in consumer-grade devices for live audio applications.
Conditioning techniques and performances should be further investigated in future work. The current conditioning blocks rely on affine transformation. Still, devices such as distortion or saturation could benefit from more complex transformations to account for the possible nonlinear influence of the parameters.